When a 3.88 g sample of solid ammonium nitrate dissolves in 60.0 g of water in a coffee cup calorimeter, the temperature drops from 23.0°C to 18.4°C. Calculate Δ

Δ

H (in kJ/mol ammonium nitrate) for the dissolution process. Assume that the specific heat of the solution is the same as that of pure water.Hints 1:

  1. Calculate q of the calorimeter using the information about the water.
  2. Determine q of the reaction from q of the calorimeter.
  3. Divide by moles of ammonium nitrate dissolved.

[Ans: 25 kJ/mol]




1
Expert's answer
2021-03-09T06:02:03-0500

Molar mass of NH4NO3 = 14+4+14+48 = 80 g/mol

So, 3.88 g is equal to

3.88/80 = 0.0485 mol

Now, mass of soln = 60+3.88 g = 63.88 g

Heat absorbed

= 63.88x4.18x(23-18.4) = 1228 J

So, 0.0485 mol absorbs 1228 J

1 mol

absorbs 25319 J

So, heat of dissolution of NH4NO3 is 25.319

kJ/mol